two planes leave Los Angeles at the same time. One head south to San Diego, while the other heads north to San Francisco. The San Diego plane flies 50mph slower than the San Francisco plane. In 30 minutes, the planes are 275 miles apart. What are their speeds
Answer by ikleyn(53763)

.

    30 minutes = 0.5 of an hour.


The total distance equation is


    0.5x + 0.5(x+50) = 275


Simplify and solve


    x + 25 = 275

    x      = 275 - 25 = 250.


ANSWER.  250 mph for the slower plane (the San Diego plane).  250 + 25 = 275 mph for the faster plane (San Francisco).

Solved.

----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.